LML Observations 140814

Isaac arrived at LML at 9:30 am. The doors to the collector had been deployed and there was a sample present in the collection jar. the sample was approximately 25ml and had a light pale yellow coloration. the sample was removed and placed in a labeled bag. the collector was then disassembled and its components were washed with 409 followed by DI water. After the collector had been cleaned a blank was taken. the blank sample measured a volume of approximately 200ml. After the blank was taken the system was reset for the next fog event.

Leave a Reply

Your email address will not be published. Required fields are marked *